A calorimeter of negligible heat capacity contains 1.00 kg of water at 303 K and 50.0 g of ice at 273 K. (a) Find the final temperature T. (b) Find the final temperature T if the mass of ice is 500 g
Answer #1

1000(30-t) = 50 * 79.7 + 50 ( t - 0)

30,000 - 1000t = 3985 + 50 t

26015 = 1050 t

t = 24.78 degree = 297.78 K = 298 K

part b )

heat require to melt 0.500 kg ice = 0.5 kg * 333.5 * kJ/kg = 167 kJ

Q available = 1 * 4.184 kJ/kg ( 0-30) = -125 kJ

availabe heat is not enough to melt ice do final temprature will be 273 K